# Residual Connections Harm Self-Supervised Abstract Feature Learning

### v1

@@ Abstract
We demonstrate that adding a weighting factor to decay the strength of identity shortcuts within residual networks substantially improves semantic feature learning in the state-of-the-art self-supervised masked autoencoding (MAE) paradigm. Our modification to the identity shortcuts within a VIT-B/16 backbone of an MAE boosts linear probing accuracy on ImageNet from 67.3% to 72.3%. This significant gap suggests that, while residual connection structure serves an essential role in facilitating gradient propagation, it may have a harmful side effect of reducing capacity for abstract learning by virtue of injecting an echo of shallower representations into deeper layers. We ameliorate this downside via a fixed formula for monotonically decreasing the contribution of identity connections as layer depth increases. Our design promotes the gradual development of feature abstractions, without impacting network trainability. Analyzing the representations learned by our modified residual networks, we find correlation between low effective feature rank and downstream task performance.
@@ 1 Introduction
Section 4 shows that implementing our novel design within an MAE yields a substantial performance boost in linear probing on ImageNet-1K [15] (72.3% from a baseline 67.3%). Ablation studies on ImageNet-100 show that smaller models equipped with our decayed identity shortcuts can outperform larger ones equipped with standard residual connections. In particular, a VIT-S/16 model [17] with our design outperforms a baseline VIT-B/16 (78.5% vs. 76.5%).
A parallel motivation for our design stems from Huh et al. [31], who show that features from residual blocks have higher rank than those produced by comparative feed-forward blocks. The smooth transition between residual and feed-forward behavior induced by our decay scheme should regularize deeper features toward exhibiting low-rank characteristics. Section 5 experimentally explores the correlation between our decayed identity shortcuts and low-rank feature representations. In the self-supervised MAE setting, as well as a supervised learning context, we validate that our method not only improves classification accuracy but also yields low-rank features with distinct cluster structure.
@@ 3.1 Decayed Identity Shortcuts
Feed-forward layers. Consider a neural network of $L$ layers. For each layer $l$ parameterized with ${\bm{\theta}}_{l}$, the operation of a feedforward neural network can be described as
@@
where ${\bm{x}}_{l}\in\mathbb{R}^{d}$ represents the output from the preceding layer, and $f_{{\bm{\theta}}_{l}}$ denotes the transformation applied at the current layer. Although it possesses a notable capability for feature abstraction, it is widely recognized that in the context of building deeper models, pure feedforward architectures are prone to vanishing gradients.
Residual connections. To address the optimization problem of vanishing gradients in deeper neural networks, ResNets [24] constructs each layer as a residual function, resulting in a modification to Eq. 1:
@@
Decayed identity shortcuts for unsupervised representation learning. Setting aside the optimization benefits brought by residual connections, we rethink the role of the residual connections from the viewpoint of representation learning. Abstraction can be viewed as invariance to local changes of input and is crucial to the disentanglement of the feature space [7]. Prior work suggests that a shortcut path of residual connections tends to preserve high-frequency fine-grained information of input [22] resulting in decreased feature abstraction. We hypothesize that this lack of abstraction harms the capability of the model to learn meaningful low-level features and that ensuring an abstract structure in the deeper layers of the neural network will help improve representation learning, especially for unsupervised tasks that often use indirect proxy objectives, such as pixel-wise reconstruction loss. Motivated by this hypothesis, we propose to downweight the contribution from the shortcut path:
@@
where $\alpha_{l}\in[0,1]$ is a rescaling factor to the residual path, controlling the information flow through the skip connection. Fully expanding this relation for a network with $L$ layers indexed from 0 to $L-1$, we have that
@@ 3.2 Implementation Strategy
Stabilizing training with residual zero initialization. The model exhibits rapid feature norm growth at the beginning of training for $\alpha_{\rm min}\leq 0.7$. We suspect that the model learns to amplify the output feature norm of $f_{{\bm{\theta}}_{l}}({\bm{x}})$ to counteract the significant decay applied to the residual connection. This growth leads to training instability and negatively impacts training convergence. To address this issue, we follow previous works [27] and initialize the weights of the final output layer in each $f_{{\bm{\theta}}_{l}}$ to zero instead of using the original Xavier uniform initialization [20]. This approach significantly enhances training stability by limiting the rate of feature norm growth and enables us to explore training with even lower values of $\alpha_{\rm min}$.
@@ 4 Large-scale Experiments
Masked Autoencoders. We implement our method within a Masked Autoencoder (MAE) [26] by replacing the residual connections in the encoder’s MLP and attention blocks with decayed identity shortcuts. The MAE operates by accepting images with a random subset of pixels masked out and learning to recover the discarded pixels. As Section 3.2 describes, we add skip connections between the encoder and decoder to facilitate learning abstract representations at the bottleneck. Since the original MAE has twice the number of encoder layers as decoder layers, we inject output from every other encoder layer into the corresponding decoder layer. To match spatial dimensions, injected encoder features are combined with learnable masked tokens before channel-wise concatenation.
@@ 4.1 ImageNet-1K
MAE[26] VIT-B/16 Recon. Loss Affine+Mask 67.3*
Ours ($\alpha_{\rm min}$ = 0.6) VIT-B/16 Recon. Loss Affine+Mask 72.3
Table 1: Accuracy of linear classifier based on pre-trained learned representations on the ImageNet-1K dataset. We list the training objective and data augmentation process of each method. Training objectives include contrastive losses, feature losses, and image reconstruction. Affine data augmentation represents random flip, random crop, and resize. Full augmentation combines affine augmentation and color distortion. With our simple modification, we substantially improve the MAE baseline by encouraging the model to learn an increased abstract features over depth. We achieve competitive performance compared to I-JEPA which uses an explicit feature loss. *We report the results using our evaluation script, while the official MAE script gives 67.8%.
@@
4.2 Qualitative Experiments.
We qualitatively evaluate the capability of our method to reconstruct test images and learn semantic image embeddings on ImageNet-100.
@@
Reconstruction. We qualitatively evaluate test images reconstructed by an MAE using our framework and images reconstructed by the original MAE. We show the reconstructed images in Figure 2. While the focus of our work is entirely to improve the representations learned by an encoder, we observe that our framework does not harm the reconstructions. Hence, there is no qualitative tradeoff for our increase in linear probing accuracy.
Embedding visualization. We also qualitatively evaluate the feature learning in Figure 3 by visualizing the t-SNE embeddings of the learned features for a random subset of test images. While the Residual connection (baseline) embeddings in Figure 3(a) produce some separation, they struggle to differentiate some harder categories, such as tubs (brown), stairs (yellow), and church buildings (purple). In contrast, the embeddings produced from our features in Figure 3(b) display much clearer separation and are able to separate almost all categories distinctly.
@@ 4.3 Ablation
Maximum decay rate. The only parameter of our framework is $\alpha_{\rm min}$, the maximum rate of decay of the identity shortcut in a residual connection. In Table 2, we show linear probing scores for varying values of $\alpha_{\rm min}$. We observe that $\alpha_{\rm min}$ must be sufficiently small to regularize the flow of information through the residual connection effectively. A reasonably small $\alpha_{\rm min}$ prevents the deeper layers of the encoder from relying heavily on the residual connections, allowing for more abstract representations in the bottleneck. This yields up to a 7.1% improvement over the ViT-B/16 baseline (${\alpha_{\rm min}}=1$) in linear probing accuracy on ImageNet-100. On the other hand, if $\alpha_{\rm min}$ is too small, for example, $\alpha_{\rm min}\leq 0.4$ for ViT-B/16, we observe that the training becomes unstable.
@@
We also explore a different parameterization of the residual connection. A few existing works leverage the benefits of re-scaling the residual connections in the context of image generation tasks [34, 57, 35, 37], but all of them scale both branches of the residual block jointly. Following this design, we try applying a constant factor $\alpha=\sqrt{0.5}$ to both ${\bm{x}}$ and $f_{{\bm{\theta}}_{l}}({\bm{x}})$, as seen in the implementation of Karras et al. [37]. In Table LABEL:tab:decay_schema, we observe that including such constant re-scaling to both branches decreases the linear probing score of the model. We also try applying a constant scaling factor to only $f_{{\bm{\theta}}_{l}}({\bm{x}})$, a case studied by [18]. These two ablations confirm that the key to abstract representation learning is controlling the weight of the identity shortcut path, leveraging the structural bias reminiscent of feedforward neural networks.
@@
Configurations Decay Block $\alpha_{\rm{min}}$ Accuracy ${\bm{x}}_{l+1}={\bm{x}}_{l}+\sqrt{0.5}f_{{\bm{\theta}}_{l}}({\bm{x}}_{l})$ MLP & Atten. -- 76.9 ${\bm{x}}_{l+1}=\sqrt{0.5}\left({\bm{x}}_{l}+f_{{\bm{\theta}}_{l}}({\bm{x}}_{l})\right)$ MLP & Atten. -- 82.6 ${\bm{x}}_{l}=\alpha_{l}{\bm{x}}_{l}+f_{\theta}({\bm{x}}_{l})$ Atten. 0.6 79.3 ${\bm{x}}_{l}=\alpha_{l}{\bm{x}}_{l}+f_{\theta}({\bm{x}}_{l})$ MLP 0.6 80.6 ${\bm{x}}_{l}=\alpha_{l}{\bm{x}}_{l}+f_{\theta}({\bm{x}}_{l})$ MLP & Atten. 0.6 83.6*
@@
5 Discussion
In this section, we try to answer a key question: How and why do residual connections impact the abstraction level of the deeper layers in a neural network? We delve deeper into how our design reinforces the low-rank bias of neural networks and try to connect our method to ideas in existing works [31]. To this end, we visualize the training dynamics of our method in both the supervised and unsupervised settings to provide a holistic analysis.
@@ 5 Discussion
Effective rank. To quantify the rank of the learned representation, Huh et al. [31] use effective rank, which is a continuous measure. For a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, the effective rank $\rho({\bm{A}})$ is defined as the Shannon entropy of the normalized singular values [51]:
@@
5.1 Abstraction and Low-rank in the Supervised Setting
(b) Effective rank of supervised ResNet over training epoch.
Inspired by their analysis, we conjecture that the improvement to feature learning capability of our method can mainly be attributed to the decayed identity shortcuts promoting low-rank features at the bottleneck. To verify our conjecture and evaluate how our method affects the effective rank of the output feature matrices, we apply our method to a supervised classification task on a 10-class subset of ImageNet, ensuring that our model is over-parameterized, following the experimental setup in [31].
We first follow the evaluation setup in [31] to analyze the impact of the model depth on feature rank when applying our method. In this experiment, we modify the standard ResNet-18 model to experiment with different depth models. By default, the ResNet-18 has a total of 8 residual blocks that are equally distributed into 4 layers. To increase model depth, we repeat residual blocks in the 3rd layer to obtain models varying between 8 and 16 total layers. At convergence, we observe that the models of different depths achieve a similar test accuracy. However, despite similar accuracies, in Figure 5(a), which visualizes the effective rank over depth for different values of $\alpha_{\rm min}$, we see that the effective rank decreases over depth. Furthermore, smaller values of $\alpha_{min}$ consistently lead to features with lower effective rank.
@@ 5.1 Abstraction and Low-rank in the Supervised Setting
5.2 Abstraction and Low-rank in the Unsupervised Setting
We would like to verify if our previous observations in supervised tasks hold for unsupervised feature learning. Specifically, we measure the training dynamics of the models trained in Table 2 (MAEs trained on ImageNet-100) in terms of accuracy and the effective rank. In Figures 6(a) and 6(b), we visualize the effective rank and linear probing accuracy of these models throughout training.
In Figure 6(a), we show that in the first 300 epochs of training, our framework can substantially decrease the effective rank of $\rho({\bm{A}}_{\bm{\theta}})$. At this stage, from Figure 6(b), we can see that our model with $\alpha_{\rm min}=0.6$ achieves a maximal improvement of 7.8% over the baseline with $\alpha_{\rm min}=1.0$. As training continues, models with lower $\alpha_{\rm min}$ consistently achieve higher probing scores. However, while the $\alpha_{\rm min}=0.5$ and 0.6 models have the lowest effective feature ranks, supporting the conjecture that low-rank features lead to better abstractions, we do observe outliers in $\alpha_{\rm min}=0.7$ and 0.8 having larger feature ranks.
Overall, along with the empirical observations found in Section 5.1 in the supervised scenario, we have uncovered evidence supporting our conjecture that an appropriate low-rank bias helps abstract representation learning and shortcut decay in residual connections promotes this bias, enhancing representation learning performance.
@@ 6 Conclusion
7 Acknowledgement
RJ and RW gratefully acknowledge the support of AFOSR grants FA9550-18-1-0166, NSF grants DMS-2023109, and NSF grants PHY2317138.

### v2

# Residual Connections Harm Abstract Feature Learning in Masked Autoencoders

@@ Abstract
We demonstrate that adding a weighting factor to decay the strength of identity shortcuts within residual networks substantially improves semantic feature learning in the state-of-the-art self-supervised masked autoencoding (MAE) paradigm. Our modification to the identity shortcuts within a VIT-B/16 backbone of an MAE boosts linear probing accuracy on ImageNet from 67.8% to 72.7%. This significant gap suggests that, while residual connection structure serves an essential role in facilitating gradient propagation, it may have a harmful side effect of reducing capacity for abstract learning by virtue of injecting an echo of shallower representations into deeper layers. We ameliorate this downside via a fixed formula for monotonically decreasing the contribution of identity connections as layer depth increases. Our design promotes the gradual development of feature abstractions, without impacting network trainability. Analyzing the representations learned by our modified residual networks, we find correlation between low effective feature rank and downstream task performance.
@@ 1 Introduction
A parallel motivation for our design stems from Huh et al. [28], who show that features from residual blocks have higher rank than those produced by comparative feed-forward blocks. The smooth transition between residual and feed-forward behavior induced by our decay scheme should regularize deeper features toward exhibiting low-rank characteristics. Section 5 experimentally explores the correlation between our decayed identity shortcuts and low-rank feature representations. Our contributions are as following:
  • $\bullet$
    We introduce decayed identity shortcuts, a novel simplified mechanism governed by a single hyper-parameter, which enhances semantic feature abstraction in masked autoencoders.
  • $\bullet$
    We identify a key correlation between our decayed identity shortcuts and low-rank inductive bias, empirically validating that our method improves classification accuracy and yields low-rank features with distinct cluster structures.
  • $\bullet$
    Our novel design within an MAE yields a substantial performance boost in linear probing on ImageNet-1K [3] (72.7% from a baseline 67.8%).
  • $\bullet$
    Ablation studies on ImageNet-100 show that smaller models equipped with our decayed identity shortcuts can outperform larger ones equipped with standard residual connections. In particular, a VIT-S/16 model [29] with our design outperforms a baseline VIT-B/16 (78.5% vs. 76.5%).
@@ 3.1 Decayed Identity Shortcuts
Feed-forward layers. Consider a neural network of $L$ layers. For each layer $l$ parameterized with ${\bm{\theta}}_{l}$, the operation of a feed-forward neural network can be described as:
@@
where ${\bm{x}}_{l}\in\mathbb{R}^{d}$ represents the output from the preceding layer, and $f_{{\bm{\theta}}_{l}}$ denotes the transformation applied at the current layer. Although it is widely known that pure feed-forward architectures are susceptible to vanishing gradients when building deeper models, Huh et al. [28] demonstrates that feed-forward modules offer implicit structural regularization, enabling deep models to generate abstract representations at bottlenecks.
Residual connections. To address the optimization problem of vanishing gradients in deeper neural networks, ResNets [2] construct each layer as a residual function, resulting in a modification to Eq. 1:
@@
Decayed identity shortcuts for unsupervised representation learning. Setting aside the optimization benefits brought by residual connections, we rethink the role of the residual connections from the viewpoint of representation learning. Abstraction can be viewed as invariance to local changes of input and is crucial to the disentanglement of the feature space [55]. Prior work suggests that a shortcut path of residual connections tends to preserve high-frequency fine-grained input information [7], resulting in decreased feature abstraction. We hypothesize that this lack of abstraction harms the capability of the model to learn meaningful low-level features and that ensuring an abstract structure in the deeper layers of the neural network will help improve representation learning, especially for unsupervised tasks that often use indirect proxy objectives, such as pixel-wise reconstruction loss. Motivated by this hypothesis, we propose to downweight the contribution from the shortcut path:
@@
where $\alpha_{l}\in[0,1]$ is a rescaling factor to the residual path, controlling the information flow through the skip connection. Fully expanding this relation for a network with $L$ layers indexed from 0 to $L-1$, we have that:
@@ 3.2 Implementation Strategy
Stabilizing training with residual zero initialization. The model exhibits rapid feature norm growth at the beginning of training for $\alpha_{\rm min}\leq 0.7$. We suspect that the model learns to amplify the output feature norm of $f_{{\bm{\theta}}_{l}}({\bm{x}})$ to counteract the significant decay applied to the residual connection. This growth leads to training instability and negatively impacts training convergence. To address this issue, we follow the implementation of previous works [19] and initialize the weights of the final output layer in each $f_{{\bm{\theta}}_{l}}$ to zero instead of using the original Xavier uniform initialization [56]. This approach significantly enhances training stability by limiting the rate of feature norm growth and enables us to explore training with even lower values of $\alpha_{\rm min}$.
@@ 4 Large-scale Experiments
Masked Autoencoders. We implement our method within a Masked Autoencoder (MAE) [1] by replacing the residual connections in the encoder’s MLP and attention blocks with decayed identity shortcuts. The MAE operates by accepting images with a random subset of pixels masked out and learning to recover the discarded pixels. As Section 3.2 describes, we add skip connections between the encoder and decoder to facilitate learning abstract representations at the bottleneck. Since the original MAE has twice the number of encoder layers as decoder layers, we inject output from every other encoder layer into the corresponding decoder layer. To match spatial dimensions, injected encoder features are combined with learnable masked tokens before channel-wise concatenation. The implementation details for the training and evaluation are shown in Section A.
@@ 4.1 ImageNet-1K
MAE[1] VIT-B/16 Recon. Loss Affine+Mask 67.8
Ours ($\alpha_{\rm min}$ = 0.6) VIT-B/16 Recon. Loss Affine+Mask 72.7
Table 1: Accuracy of linear classifier based on pre-trained learned representations on the ImageNet-1K dataset. We list the training objective and data augmentation process of each method. Training objectives include contrastive losses, feature losses, and image reconstruction. Affine data augmentation represents random flip, random crop, and resize. Full augmentation combines affine augmentation and color distortion. With our simple modification, we substantially improve the MAE baseline by encouraging the model to learn an increased abstract features over depth. We achieve competitive performance compared to I-JEPA which uses an explicit feature loss.
@@
Embedding visualization. We qualitatively evaluate the feature learning in Figure 2 by visualizing the t-SNE embeddings of the learned features for a random subset of test images. While the Residual connection (baseline) embeddings in Figure 2(a) produce some separation, they struggle to differentiate some harder categories, such as tubs (brown), stairs (yellow), and church buildings (purple). In contrast, the embeddings produced from our features in Figure 2(b) display much clearer separation and are able to separate almost all categories distinctly.
@@ 4.2 Ablation
Decay rate $\alpha_{\rm min}$. The only parameter of our framework is $\alpha_{\rm min}$, the minimum scaling factor applied to the identity shortcut at the final layer. In Table 2,we show linear probing scores for varying values of $\alpha_{\rm min}$. We observe that $\alpha_{\rm min}$ must be sufficiently small to regularize the flow of information through the residual connection effectively. A reasonably small $\alpha_{\rm min}$ prevents the deeper layers of the encoder from relying heavily on the residual connections, allowing for more abstract representations in the bottleneck. This yields up to a 7.1% improvement over the ViT-B/16 baseline (${\alpha_{\rm min}}=1$) in linear probing accuracy on ImageNet-100. On the other hand, if $\alpha_{\rm min}$ is too small, for example, $\alpha_{\rm min}\leq 0.4$ for ViT-B/16, we observe that the training becomes unstable.
@@
We also explore a different parameterization of the residual connection. A few existing works leverage the benefits of re-scaling the residual connections in the context of image generation tasks [62, 63, 64, 65], but all of them scale both branches of the residual block jointly. Following this design, we try applying a constant factor $\alpha=\sqrt{0.5}$ to both ${\bm{x}}$ and $f_{{\bm{\theta}}_{l}}({\bm{x}})$, as seen in the implementation of Karras et al. [65]. In Table LABEL:tab:decay_schema, we observe that including such constant re-scaling to both branches decreases the linear probing score of the model. We also try applying a constant scaling factor to only $f_{{\bm{\theta}}_{l}}({\bm{x}})$, a case studied by [27]. Decaying $f_{{\bm{\theta}}_{l}}$ term does not alter the results much since the model can easily learn a scaling factor within $f_{{\bm{\theta}}_{l}}$ to counterbalance the added decaying force. These two ablations confirm that the key for abstract representation learning is to control the weights of the identity shortcuts, leveraging the structural bias reminiscent of feed-forward neural networks.
@@
Configurations Decay Block $\alpha_{\rm{min}}$ Accuracy ${\bm{x}}_{l+1}={\bm{x}}_{l}+f_{{\bm{\theta}}_{l}}({\bm{x}}_{l})$- 76.5 ${\bm{x}}_{l+1}={\bm{x}}_{l}+\sqrt{0.5}f_{{\bm{\theta}}_{l}}({\bm{x}}_{l})$ MLP & Atten. -- 76.9 ${\bm{x}}_{l+1}=\sqrt{0.5}\left({\bm{x}}_{l}+f_{{\bm{\theta}}_{l}}({\bm{x}}_{l})\right)$ MLP & Atten. -- 82.6 ${\bm{x}}_{l}=\alpha_{l}{\bm{x}}_{l}+f_{\theta}({\bm{x}}_{l})$ Atten. 0.6 79.3 ${\bm{x}}_{l}=\alpha_{l}{\bm{x}}_{l}+f_{\theta}({\bm{x}}_{l})$ MLP 0.6 80.6 ${\bm{x}}_{l}=\alpha_{l}{\bm{x}}_{l}+f_{\theta}({\bm{x}}_{l})$ MLP & Atten. 0.6 83.6*
@@
5 Discussion on feature rank
In this section, we try to answer a key question: How and why do residual connections impact the abstraction level of the deeper layers in a neural network? We delve deeper into how our design reinforces the low-rank bias of neural networks and try to connect our method to ideas in existing works [28]. To this end, we visualize the training dynamics of our method and analysis the feature rank of our approach to provide a holistic analysis.
@@ 5 Discussion on feature rank
Effective rank. For analysis purpose, Huh et al. [28] quantify the rank of the learned representation using the effective rank, which is a continuous measure. For a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, the effective rank $\rho({\bm{A}})$ is defined as the Shannon entropy of the normalized singular values [66]:
@@
Inspired by their analysis, we conjecture that the improvement to feature learning capability of our method can mainly be attributed to the decayed identity shortcuts promoting low-rank features at the bottleneck. We measure the training dynamics of the models presented in Table 2 (MAEs trained on ImageNet-100) in terms of accuracy and the effective rank.
In Figure 4(a) and 4(b), we present the training dynamics of our model, illustrating effective rank and linear probing accuracy across various values of $\alpha_{\text{min}}$. During training, models with lower $\alpha_{\text{min}}$ tend to exhibit higher accuracy and lower effective rank, which supports the conjecture that low-rank features contribute to better learned representations in the MAE model. Additionally, we observe that these patterns persist throughout the training process, suggesting that establishing a low-rank bottleneck from the outset benefits model training right from the beginning.
Overall, we have uncovered evidence supporting our conjecture that an appropriate low-rank bias helps abstract representation learning and shortcut decay in residual connections promotes this bias, enhancing the abstract feature learning performance.
Compatibility with contrastive learning frameworks. Though seeing the big improvement of applying our decayed identity shortcuts in masked autoencoders, we note that our approach does not easily extend to contrastive learning frameworks, where the low rank inductive bias conflicts with the training objectives. For example, the training objective in MoCov3 [58] include an universal repulsive term in the denominator, which is supposed to increase the feature rank. Furthermore, contrastive learning also imposes specific data distribution requirements, such as both image crops needing to include the same object, which is impractical for in-the-wild images. In contrast, masked autoencoders do not have this requirement, offering greater potential for scalability and generalization. Our work focuses on further improving the learned representations of masked autoencoders, and enhancing contrastive learning is beyond the scope of our discussion.
@@ 6 Conclusion
Appendix A Training and Evaluation Details
A.1 Model training.
Our training configurations primarily followed the guidelines established by He et al. [1]. In the ImageNet-1K experiment, our model was trained for 800 epochs, utilizing the AdamW [67] optimizer with a constant weight decay of 5e-2 for a batch size of 1024. We set the maximum learning rate to 6e-4. Initially, the learning rate started at 0 and linearly increased to its maximum over the first 40 epochs, after which it followed a cosine schedule to gradually decrease to zero by the end of the training period. It is worth noting that the learning rate per sample, or effective learning rate, in our setup matched that of He et al. [1], although our maximum learning rate was set lower due to our batch size being a quarter of theirs. We applied random resizing, cropping, and horizontal flipping during training as part of our augmentation scheme. To enhance the quality of the learned representations in most experiments, we employed the normalized pixel loss, as proposed by [1]. For visualization purposes, as shown in Figure 6, we used L2 pixel-wise loss. In the ImageNet-100 experiment, we employed the identical training configuration used in the ImageNet-1K experiments. We train our model with 4 NVIDIA A40 GPUs and a completed trianing usually takes 20 hours on ImageNet-100 and 200 hours on ImageNet-1k.
A.2 Evaluation with Linear Probing.
For the ImageNet-1k dataset, we use the exact same evaluation protocols employed in He et al. [1], which includes random data augmentation.
For the ImageNet-100 dataset, we employed a simpler evaluation protocol: We train the linear classifier with a batch size of 1024 for 200 epochs, where the learning rate starts at 1e-2 and then decays towards 0 using a cosine scheduler. During this evaluation, we do not apply any data augmentation.
Figure 5: We present our enhanced UNet Transformer architecture for Masked Auto-encoder. (1) Left: Our customized encoder blocks, equipped with our proposed decay identity shortcuts. (2) Middle: Standard transformer blocks as the decoder blocks. (3) Right: We incorporate the decay identity shortcuts exclusively within the encoder blocks of our UNet transformer and employ standard transformer blocks for the decoder. To support abstract representation learning at the bottleneck, i.e., the last layer of the Encoder 12, we adopt the UNet [25] architecture and create skip connections that transmit every other encoder feature directly to the decoder.
A.3 Modified Architecture
We present a visualization of our UNet transformer design, as outlined in Section  3.2, in Fig. 5. It’s important to note that decayed identity shortcuts are exclusively implemented within the encoder block. Additionally, we establish skip connections from alternating blocks in the encoder to the decoder, following the UNet [68] architecture’s design principles.
Appendix B Further experiments
B.1 Further Ablation of Maximum Decay Rate on ImageNet-1K
Dataset Backbone $\alpha_{\rm min}$ 0.5 0.6 0.7 0.8 0.9 1.0
ImageNet-1K ViT-B/16 69.8 72.7 68.9 - - -
ImageNet-100 ViT-B/16 82.3 83.6 81.8 79.8 79.2 76.5
ImageNet-100 ViT-S/16 78.6 78.5 78.1 75.2 73.5 69.2
Table 4: Linear probing accuracy of our method by varying $\alpha_{\rm min}$, the architecture size, and the dataset. We extend the ablation studies detailed in Table. 2 by including linear probing results for the ImageNet-1K dataset. The results indicate that the most favorable $\alpha_{\rm min}$ consistently falls within the range of $[0.5,0.6]$ for ImageNet-1K and ImageNet-100 experiments.
We present the results of ablating the choices of $\alpha_{\rm min}$ on ImageNet-1K dataset in Table.4. From the table, we show that the optimal $\alpha_{\rm min}$ for ViT-B/16 on ImageNet-100 matches the optimal one for ImageNet-1K, while a lower $\alpha_{\rm min}$ is preferred for a smaller architecture ViT-S/16.
B.2 Reconstruction quality.
We qualitatively evaluate test images reconstructed by an MAE using our framework and images reconstructed by the original MAE. We show the reconstructed images in Figure 6. While the focus of our work is entirely to improve the representations learned by an encoder, we observe that our framework does not harm the reconstructions. Hence, there is no qualitative tradeoff for our increase in linear probing accuracy.
B.3 Abstraction and Low-rank in the Supervised Setting
In this experiment, we modify the standard ResNet-18 model to experiment with different depth models. By default, the ResNet-18 has a total of 8 residual blocks that are equally distributed into 4 layers. To increase model depth, we repeat residual blocks in the 3rd layer to obtain models varying between 8 and 16 total layers. At convergence, we observe that the models of different depths achieve a similar test accuracy. However, despite similar accuracies, in Figure 7(a), which visualizes the effective rank over depth for different values of $\alpha_{\rm min}$, we see that the effective rank decreases over depth. Furthermore, smaller values of $\alpha_{min}$ consistently lead to features with lower effective rank.
(b) Effective rank of supervised ResNet over training epoch.